OpenStudy (ttp):

Show your work to prove that the equation y = 12x + 8 correctly represents the function for the data in this table.

OpenStudy (helder_edwin):

u have three values for x. replace them in the equation and if u get the values already provided for y, then u got the prove.

OpenStudy (helder_edwin):

replace each value of x into the equation. do the arithmetic and see if u get the corresponding value for y.

OpenStudy (helder_edwin):

for instance, if x=2 then \[\large y=12x+8=12\cdot2+8=24+8=32 \] as expected. do the same with the other two values of x.